OpenStudy (kainui):

Without expanding evaluate the determinant:

OpenStudy (kainui):

\[\LARGE \left|\begin{matrix}a^2 &a & 1 \\ b^2 & b & 1 \\ c^2 & c & 1\end{matrix}\right|\]

OpenStudy (anonymous):

Ya you told me trick , it slipped of my mind :( ,

OpenStudy (anonymous):

the*

OpenStudy (kainui):

Also, here's some extra info in case you were wondering. \[\Large a \ne b \ne c\] if any of them are equal the whole determinant is obviously 0 haha.

ganeshie8 (ganeshie8):

first observation is \[ \large \left|\begin{matrix}a^2 &a & 1 \\ b^2 & b & 1 \\ c^2 & c & 1\end{matrix}\right| = F(a,b,c) \] 3rd degree function

OpenStudy (ikram002p):

like this ? \(det \left ( \begin{bmatrix} a^2 & a & 1\\ b^2 & b& 1\\ c^2& c & 1 \end{bmatrix} \right ) = det \left ( a^2\begin{bmatrix} b& 1\\ c&1 \end{bmatrix}-a \begin{bmatrix} b^2& 1\\ c^2&1 \end{bmatrix}+ \begin{bmatrix} b^2& b\\ c^2&c \end{bmatrix} \right ) \) then what ?

OpenStudy (kainui):

Yeah you're not allowed to do that @ikram002p heh =P

OpenStudy (ikram002p):

ohh lol without expantion hehe i thought with expantion and said wth xD

OpenStudy (ikram002p):

well in case c= n (a) b=m (a) this might be weird xD

ganeshie8 (ganeshie8):

\(a=b\) makes \(F(a,b,c) = 0\) so \((a-b)\) is a factor of \(F(a,b,c)\) : \[F(a,b,c) = (a-b)(something)\]

OpenStudy (kainui):

That's on the right path I think ganeshie! I got to that result by a different method though so I am really surprised, very cool!

OpenStudy (ikram002p):

wait :o how this matrix relatied to this \(\begin{bmatrix} 1& 1 &1 \\ a&b &c \\ a^2& b^2 & c^2 \end{bmatrix} \) ?

OpenStudy (kainui):

I dunno @ikram002p lol. That's like the backwards transpose (is that a thing?) idk lol

OpenStudy (ikram002p):

lol i dont know as well :P

ganeshie8 (ganeshie8):

Kainui, il conclude : similar reasoning yields (b-c) and (c-a) as other facors, we stop at this point cuz it is clear that F is a 3rd degree function So F(a,b,c) = k(a-b)(b-c)(c-a)

ganeshie8 (ganeshie8):

k has to be 1 for obvious reasons. cant wait to see the another method of solving this xD

OpenStudy (kainui):

You almost got it, but I think you're wrong by a negative sign! What I did was subtract the second row from the first and factor out a (a-b) d=k(a-b)(b-c)(c-a) now we know that the diagonal is one of the elements of the determinant, a^2b so if we look in our determinant we have -ka^2b so then k=-1.

OpenStudy (ikram002p):

:o

OpenStudy (kainui):

\[\LARGE \left|\begin{matrix}a^2-b^2 & a-b & 0\\ b^2 & b &1 \\ c^2 & c &1\end{matrix}\right|\]\[\LARGE (a-b)\left|\begin{matrix}a+b & 1 & 0\\ b^2 & b &1 \\ c^2 & c &1\end{matrix}\right|\]

OpenStudy (xapproachesinfinity):

@bradely you killed the fun my friend lol following this guys how they were approaching the problem was interesting

OpenStudy (ikram002p):

ill sighn in in that site :P